### v1

@@
The Cost of Coming Outthanks:  We especially would like to thank Michael Lechner and Franco Peracchi for feedback and suggestions. We are also grateful to Jaime Arellano-Bover, Nora Bearth, Jonathan Chassot, Caroline Coly, Daniel Goller, Giorgio Gulino, Moritz Janas, David Neumark, Paolo Pinotti, Mounu Prem, Dario Sansone, Erik-Jan Senn, seminar participants at University of Rome Tor Vergata and SEW-HSG research seminars, and conference participants at the 2nd Rome Ph.D. in Economics and Finance Conference for comments and discussions.
Enzo Brox  Swiss Institute for Empirical Economic Research (SEW), University of St.Gallen.    Riccardo Di Francesco  Department of Economics and Finance, University of Rome Tor Vergata, Rome. Electronic correspondence: riccardo.di.francesco@uniroma2.it.
(March 6, 2024)
The fear of social stigma and discrimination leads many individuals worldwide to hesitate in openly disclosing their sexual orientation. Due to the large costs of concealing identity, it is crucial to understand the extent of anti-LGB sentiments and reactions to coming out. However, disclosing one’s sexual orientation is a personal choice, complicating data access and introducing endogeneity issues. This paper tackles these challenges by using an innovative data source from a popular online video game together with a natural experiment. We exploit exogenous variation in the identity of a playable character to identify the effects of disclosure on players’ revealed preferences for that character. Leveraging detailed daily data, we monitor players’ preferences for the character across diverse regions globally and employ synthetic control methods to isolate the effect of the disclosure on players’ preferences. Our findings reveal a substantial and persistent negative impact of coming out. To strengthen the plausibility of social stigma as the primary explanation for the estimated effects, we systematically address and eliminate several alternative game-related channels.
@@ 1 Introduction
In this paper, we use an innovative data source from an online video game together with a natural experiment to assess the stigma attached to sexual minority status by observing individuals’ responses to sexual minority disclosure. Existing studies examining anti-LGB sentiments, which we review below, usually face two limitations. First, they rely on a selection-on-observables identification strategy that is likely to suffer from omitted variable bias. Second, they depend on survey data where individuals have the discretion to choose whether to disclose their sexual orientation, thereby introducing additional endogeneity issues [168, 206]. Our use of video game data effectively avoids these selection concerns, and the natural experiment we leverage allows for a credible identification of the responses to revealing sexual minority status.
We use a rich data set from one of the most popular online video games, League of Legends. In this game, before a match starts, players are required to choose one playable character. Each character is characterized by game-relevant attributes and a background story that provides details about their history, origin, and relationships with other in-game characters. We leverage an unexpected change in the background story of a playable character, which discloses its sexual orientation minority status, to examine individuals’ responses to sexual minority disclosure.
@@
Our findings reveal a substantial and persistent negative impact of coming out on players’ preferences for the treated character, with a decline of more than $30\%$ of the pre-treatment average preferences for that character. This result consistently holds across various robustness checks. Additionally, we exploit another unique feature of our setting. The online video game is played by individuals around the world under very comparable circumstances but on different regional servers. We make use of information on the regional servers to compare how preferences for the treated character evolve across diverse regions worldwide. The results consistently demonstrate a negative response across regions.
To strengthen the credibility of preferences towards LGB status as the primary explanation for the estimated effects, it is crucial to ensure that players’ decisions to switch from the character are not influenced by other factors. We address and eliminate several alternative channels, thereby enhancing the plausibility of social stigma as the primary explanation for the observed behavior. First, we rule out the possibility that shifts in characters’ relative strengths could explain our estimated effect. Second, we show that players’ skills have no correlation with the choice to drop the character, thus dismissing the possibility that gameplay factors are the driving force behind the players’ observed behavior. Additionally, we demonstrate that players are not leaving the game after the disclosure but are shifting their focus to other characters. Third, we provide evidence that switching to other characters does not affect the performance of the players involved, highlighting that the decision to abandon the character is not driven by performance considerations.
@@
While our setting is unconventional, it presents unique advantages for the identification of sentiments towards LGB identity. First, video games offer a controlled research environment, enabling the observation of behaviors that might be challenging to capture through traditional methods [202].222 The recognition of video game data’s potential for research is growing among economists, although at a gradual pace. [173], [203], and [200] are instances of this emerging trend. [203] investigate consumer behavior in the video game industry. They examine the impact of marking products with a gay label on consumers’ demand, finding a significant, albeit short-lived, decrease in demand following the introduction of the gay label. Second, they allow us to leverage objective measures of behavior and identity, circumventing the limitations associated with self-reported identity in surveys [168, 206]. Conventional survey questions have been shown to result in an underrepresentation of the LGB community, thus introducing additional selection concerns. Third, online gaming platforms offer the benefit of anonymity, minimizing social desirability bias and increasing the likelihood of individuals disclosing sensitive information such as their true attitudes toward sexual minority groups.
1.1 Relation to literature
@@
Our paper contributes to three distinct strands of the literature. First, it relates to a growing literature studying the economics of LGBTQ+ individuals [[, see, e.g.,]]badgett2021lgbtq, badgett2023review. The current body of research primarily focuses on measuring discrimination against LGB individuals by comparing their labor market outcomes with those of non-minority individuals with similar observable characteristics, mostly documenting wage penalties for gay and bisexual men and wage premiums for lesbian women [[, see, e.g.,]]badgett1995wage, martell2021labor, carpenter2017does, carpenter2023orient.333 The only study finding a wage premium for gay men is that of [167]. [207] does not find a wage penalty but finds a lower probability of being full-time employed and a higher probability of being a victim of work-related violence for gay men. There is also evidence that transgender workers face earning and employment penalties [170, 191]. However, despite their valuable contributions, these studies suffer from the endogeneity and selection issues discussed above that hinder the ability to draw causal inferences from their findings.444 A related literature studies whether changes in laws and norms affect labor market outcomes and attitudes towards LGB individuals [[, see, e.g.,]]burn2018not, sansone2019pink, ofosu2019same, burn2020relationship, delhommer2020effect, deal2022bound, deal2023heterogeneity. [164] show that a randomized intervention that encourages actively taking the perspective of others can reduce transphobia. To tilt towards a more causal interpretation, other studies use correspondence designs to probe into hiring discrimination against LGB individuals, consistently revealing that LGB job candidates are less likely to be invited for interviews or offered job opportunities [148, 150, 156, 157, 159]. Nevertheless, a significant challenge lies in communicating that an individual belongs to a sexual orientation minority group, since such information is not typically included in job applications. This raises the question of whether the observed results are affected by the choice and nature of the used signal [165].
@@
We make several contributions to this literature. First, our use of data from an online video game allows us to overcome the endogeneity issues of previous studies and to identify preferences toward LGB status. Second, to the best of our knowledge, our study is the first to investigate the immediate reactions to coming out. Due to the substantial anecdotal and scientific evidence that individuals often hide their sexual orientation status [[, see, e.g.,]]badgett2020economic and the prevalent evidence of the associated costs [147, 181], several studies have explicitly focused on investigating the determinants and incentives of coming-out decisions. [188] investigate the impact of same-sex marriage legalization on coming-out decisions using a revealed preference mechanism inferred from data on Catholic priests’ vow of celibacy, finding reduced demand for priestly studies after the adoption of same-sex legalization. [193] explore spillover effects of coming-out decisions using Twitter data, discovering positive externalities, as exposure to peers coming out is associated with a higher probability of individuals coming out themselves. [196] conduct a lab experiment demonstrating that individuals strategically hide their sexual orientation in anticipation of discrimination in prosocial behavior, a result consistent with that of [194]. Our study complements these studies by focusing on responses to revealing sexual minority status, potentially providing a rationale for the observed underrepresentation of individuals with sexual minority status in many areas.
Second, our contribution extends to the literature exploring the intersection of identity and decision-making, as discussed by [145]. Empirical investigations encounter a challenge where causality is likely bidirectional - meaning that identity can be a cause or a consequence. Furthermore, identity is a multifaceted concept, thus isolating a particular component of identity is empirically challenging [182]. Previous research overcomes these challenges and documents in-group and out-group behavior using either individuals’ inherent identities [[, see, e.g.,]]giuliano2009manager,price2010racial, oh2023does or by experimentally introducing identities [[, see, e.g.,]]chen2009, chen2011.555 For recent reviews of the literature, see [179, 182, 171]. Results highlight the importance of an individual’s identity in shaping their behaviors, revealing a tendency for preferential treatment or bias towards those who share similar characteristics (in-group bias). To the best of our knowledge, our study is among the first studies to causally show the influence of sexual orientation status on preferences. Because sexual orientation is a key but non-salient facet of social identity, our paper makes an important contribution to this literature.
@@
Third, our findings also contribute to the broader literature on discrimination. Early work by [140] paved the way for an extensive literature investigating discrimination instances, mostly based on gender and ethnicity, across diverse economic domains [[, see, e.g.,]]arnold2022measuring, kuhn2013gender. A substantial portion of the empirical evidence stems from field experiments, in which researchers use audit and correspondence studies to isolate the causal impact of identity on behavior [[, see, e.g.,]]ayres1995race, oreopoulos2011.666 For recent surveys, see [165] and [172]. [195] provides a detailed review of traditional statistical and taste-based discrimination models, along with a discussion of recent theories that expand on these models. Our study contributes by exploring sentiments against sexual orientation minorities. The concealable nature of sexual orientation allows individuals to anticipate discrimination and strategically choose to hide their identity [196]. This complicates the use of traditional methods to investigate discrimination against individuals based on their sexual orientation and necessitates innovative strategies for understanding sentiments against groups with stigmatized identities, particularly in environments prone to discrimination.
@@ 2.1 League of Legends
League of Legends is a prominent multiplayer online game developed and published by Riot Games. Originally launched on October $27^{th}$, $2009$, the game attracted an impressive player base, averaging $180$ million monthly active players as of $2022$ and reaching a peak of $14$ million players in one day. League of Legends has also achieved significant financial success, generating \$$\$1.8$ billion in revenue in $2022$.777 See, e.g., https://prioridata.com/data/league-of-legends.
@@
Before a match begins, players must select a playable character to control during the match from a pool of $165$ available characters. In our analysis, we measure players’ revealed preferences for a specific character by quantifying how frequently they select that character for their matches. Our objective is to investigate whether these preferences undergo any shifts following the disclosure of the character’s sexual orientation. Thus, we devote the rest of this section to exploring the design of characters in League of Legends and the process through which players select their characters for matches.
Each character has a unique set of skills and abilities and is specifically designed to excel in one or two of the distinct roles that players can assume within the team. Additionally, characters are crafted with a rich background that adds a narrative dimension to the game but does not have any impact on the game’s mechanics. This is achieved through the creation of detailed biographies and short stories that provide players with a deeper understanding of the character’s history and motivations, thus offering players the opportunity to connect with their chosen characters on a more personal level.888 The list of all characters along with details about their abilities and histories is available at https://www.leagueoflegends.com/en-gb/champions/.
@@ 2.2 Coming-Out Event
Every year in June, LGBT Pride Month, a dedicated time to honor and celebrate the LGBT community, takes place. Since $2018$, Riot Games has actively participated in this month-long celebration by integrating new content into League of Legends during the month of June. This includes the introduction of in-game cosmetics, such as character skins, as well as emotes that allow players to express themselves in the game. It is important to note that while these additions enhance the visual and expressive elements of the game, they do not alter the game’s mechanics or the characteristics and abilities of the League of Legends characters.999 We check this in Section 5.1, where we demonstrate that characters’ strength was unaffected by LGBT Pride Month.
Riot has also supported the representation of the LGBT community in the game by unveiling the sexual minority status of specific characters.101010 For instance, in $2018$, the character Neeko was confirmed as lesbian, marking her as the first openly LGBT character in the game. Furthermore, in $2021$, the characters Diana and Leona were revealed to be lesbians, while the character Nami was declared to be lesbian and polyamorous. In our paper, we primarily focus on the character Graves, chosen due to the existence of a well-defined announcement regarding his sexual orientation.
At the beginning of the $2022$ LGBT Pride Month, Riot Games released a short story featuring Graves and Twisted Fate. This story officially unveils Graves’ sexual orientation, establishing him as a gay character.111111 The story unveiling Graves’ sexual orientation also subtly hints at Twisted Fate’s pansexuality, although this is not explicitly stated. We investigate whether this implied revelation has captured the players’ attention in Appendix C. We highlight the relatively low attention directed towards Twisted Fate from players, who were primarily focused on Graves and the explicit establishment of his sexual orientation. As a result, we concentrate our analysis on Graves and his disclosure for a more credible identification of the effects of coming out. The following quotes provide two pivotal passages of the narrative:121212 The whole story is available at https://universe.leagueoflegends.com/en_SG/story/the-boys-and-bombolini/.
@@
This coming-out event closely approximates an ideal experiment where individuals randomly disclose their sexual minority status, thus providing a unique setting to identify the effects of coming out on players’ preferences for Graves.131313 It is crucial to distinguish between the coming-out event and the disclosure of Graves’ sexual orientation. The coming-out event encompasses both Graves’ disclosure and the start of LGBT Pride Month. While this is not a concern for identification, it requires careful interpretation of the findings. To maintain clarity, we generally refer to the effects of the coming-out event in our analysis. Further discussion on this topic is deferred to Section 5.4 and Appendix D.
Figure 2.1: Google search interest over time for the queries “Graves gay” (top panel) and “LoL Graves” (bottom panel). The dashed vertical line denotes the week of disclosure, and the shaded area highlights the League of Legends World Championship.
To ensure the credibility of our identification, it is crucial that the disclosure was not anticipated by players. The top panel of Figure 2.1 displays the Google search interest for the query “Graves gay”. We observe minimal interest in this search term throughout the year $2022$, with a remarkable spike occurring during the week of the coming-out event. This pattern supports our assumption of no anticipation and strengthens the credibility of our identification strategy.
Furthermore, the lower panel of Figure 2.1 displays the Google search interest for the query “LoL Graves”.141414 LoL is the common short-hand for League of Legends. Similarly to the previous search term, we observe a remarkable spike in interest during the week of the treatment. What is particularly interesting is that this surge in interest surpasses the level observed during the $2022$ League of Legends World Championship (held from September $29^{th}$ to November $5^{th}$), despite Graves being among the top-eight most played characters during the tournament. This finding emphasizes the substantial impact and attention that the coming-out event received from players.

### v3

@@
The Cost of Coming Outthanks:  We especially would like to thank Michael Lechner and Franco Peracchi for feedback and suggestions. We are also grateful to Jaime Arellano-Bover, Nora Bearth, Jonathan Chassot, Caroline Coly, Daniel Goller, Eric Guan, Giorgio Gulino, Moritz Janas, Giovanni Mellace, David Neumark, Paolo Pinotti, Mounu Prem, Dario Sansone, Erik-Jan Senn, Andrea Smurra, seminar participants at University of Rome Tor Vergata and SEW-HSG research seminars, and conference participants at the 2nd Rome Ph.D. in Economics and Finance Conference, the SES 2024, the 29th IPDC, and the ESPE 2025 for comments and discussions.
Enzo Brox  University of Bern, Center for Research in Economics of Education and University of St.Gallen, Swiss Institute for Empirical Economic Research. Electronic correspondence: enzo.brox@unibe.ch.    Riccardo Di Francesco  University of Southern Denmark, Department of Economics, Econometrics and Data Science. Electronic correspondence: rdif@sam.sdu.dk.
(August 7, 2025)
The fear of social stigma leads many individuals worldwide to hesitate in disclosing their sexual orientation. Since concealing identity is costly, it is crucial to understand the extent of anti-LGB sentiments and reactions to coming out. This paper uses an innovative data source from a popular online game together with a natural experiment to overcome existing data and endogeneity issues. We exploit exogenous variation in the identity of a character to identify the effects of coming out on players’ revealed preferences for that character across diverse regions globally. Our findings reveal a substantial and persistent negative impact of coming out.
@@ 1 Introduction
In this paper, we use an innovative data source from an online video game together with a natural experiment to assess the stigma attached to sexual minority status by observing individuals’ responses to sexual minority disclosure. Existing studies examining anti-LGB sentiments, which we review below, usually face at least one of two limitations. First, they rely on a selection-on-observables identification strategy that is likely to suffer from omitted variable bias. Second, they depend on survey data where individuals have the discretion to choose whether to disclose their sexual orientation, thereby introducing additional endogeneity issues [165, 201]. Our use of video game data effectively avoids these selection concerns, and the natural experiment we leverage allows for a credible identification of the responses to revealing sexual minority status.
We use a rich data set from one of the most popular online video games, League of Legends. In this game, before a match starts, players are required to choose one playable character. Each character is characterized by game-relevant attributes and a background story that provides details about their history, origin, and relationships with other in-game characters. We leverage an unexpected change in the background story of a playable character, which discloses its sexual orientation minority status, to examine individuals’ responses to sexual minority status disclosure.
@@
Our findings reveal a substantial and persistent negative impact of coming out on players’ preferences for the treated character, with a decline of more than $40\%$ of the pre-treatment average preferences for that character. This result consistently holds across various robustness checks. To further assess the external validity of our findings, we conduct a heterogeneity analysis based on players’ skill levels. The decline in preferences is equally pronounced among both lower- and higher-skilled players, indicating that the behavioral response is widespread rather than confined to a particular segment of the player base. Additionally, we exploit another unique feature of our setting: the online video game is played globally on different regional servers under very comparable circumstances. We make use of the information on the regional servers to compare how preferences for the treated character evolve across diverse regions across the world. The results consistently demonstrate a negative response across regions.
To strengthen the credibility of preferences towards LGB status as the primary explanation for the estimated effects, it is crucial to ensure that players’ decisions to switch from the character are not influenced by other factors. We address and eliminate several alternative channels, thereby enhancing the plausibility of social stigma as the primary explanation for the observed behavior. First, we rule out the possibility that shifts in characters’ relative strengths could explain our estimated effect. Second, we show that players’ skills have no correlation with the choice to drop the character, thus dismissing the possibility that gameplay factors are the driving force behind the players’ observed behavior. Additionally, we demonstrate that players are not leaving the game after the disclosure but are shifting their focus to other characters. Third, we provide evidence that switching to other characters does not affect the performance of the players involved, highlighting that the decision to abandon the character is not driven by performance considerations. Fourth, we dismiss the possibility that the release of a new character after the disclosure explains the estimated effect.
@@
While our setting is unconventional, it presents unique advantages for the identification of sentiments towards LGB identity. First, video games offer a controlled research environment, enabling the observation of behaviors that might be challenging to capture through traditional methods [204].222 A similar argument underlies several studies that use data from the professional sports industry to examine, for example, discrimination [150, 170], incentive effects [152], and other behavioral patterns [203]. Second, they allow us to leverage objective measures of behavior and identity, circumventing the limitations associated with self-reported identity in surveys [165, 201]. Third, online gaming platforms offer the benefit of anonymity, minimizing social desirability bias and increasing the likelihood of individuals disclosing sensitive information such as their true attitudes toward sexual minority groups.
Despite these advantages, our study’s findings should be interpreted with three important caveats. First, the sample of participants in our setting is unlikely to be representative of the general population. The demographic composition of online gaming communities may differ systematically from broader societal distributions, potentially affecting the external validity of our findings. Second, our analysis focuses on responses to a male character revealing his sexual minority status. Given prior evidence of differential labor market discrimination against gay men and lesbian women [192], it is plausible that reactions would vary if a female character disclosed the same identity. Third, our study examines preferences for a fictitious character rather than an actual human being. As a result, factors such as the character’s perceived prestige or popularity may influence individuals’ responses, making our setting most analogous to the context of supporting a well-known public figure, such as a politician or professional athlete.
@@
Our paper contributes to three distinct strands of the literature. First, it relates to a growing literature studying the economics of LGBTQ+ individuals [[, see, e.g.,]]badgett2021lgbtq, badgett2023review. The current body of research primarily focuses on measuring discrimination against LGB individuals by comparing their labor market outcomes with those of non-minority individuals with similar observable characteristics, mostly documenting wage penalties for gay and bisexual men and wage premiums for lesbian women [[, see, e.g.,]]badgett1995wage, martell2021labor, carpenter2017does, carpenter2023orient.333 The only study finding a wage premium for gay men is that of [164]. [202] does not find a wage penalty but finds a lower probability of being full-time employed and a higher probability of being a victim of work-related violence for gay men. There is also evidence that transgender workers face earning and employment penalties [167, 186]. However, despite their valuable contributions, these studies suffer from the endogeneity and selection issues discussed above that hinder the ability to draw causal inferences from their findings.444 A related literature studies whether changes in laws and norms affect labor market outcomes and attitudes towards LGB individuals [[, see, e.g.,]]burn2018not, sansone2019pink, ofosu2019same, burn2020relationship, delhommer2020effect, deal2022bound, deal2023heterogeneity. [161] show that a randomized intervention that encourages actively taking the perspective of others can reduce transphobia. To tilt towards a more causal interpretation, other studies use correspondence designs to probe into hiring discrimination against LGB individuals, consistently revealing that LGB job candidates are less likely to be invited for interviews or offered job opportunities [147, 148, 153, 155, 157]. Nevertheless, a significant challenge lies in communicating that an individual belongs to a sexual orientation minority group, since such information is not typically included in job applications. This raises the question of whether the observed results are affected by the choice and nature of the signal used [162].
@@
We make several contributions to this literature. First, our use of data from an online video game allows us to overcome the endogeneity issues of previous studies and to identify preferences toward LGB status. Second, to the best of our knowledge, our study is the first to investigate the immediate reactions to coming out. Due to the substantial anecdotal and scientific evidence that individuals often hide their sexual orientation status [[, see, e.g.,]]badgett2020economic and the prevalent evidence of the associated costs [146, 177], several studies have explicitly focused on investigating the determinants and incentives of coming-out decisions.555 [183] investigate the impact of same-sex marriage legalization on coming-out decisions using a revealed preference mechanism inferred from data on Catholic priests’ vow of celibacy, finding reduced demand for priestly studies after the adoption of same-sex legalization. [188] explore spillover effects of coming-out decisions using Twitter data, discovering positive externalities, as exposure to peers coming out is associated with a higher probability of individuals coming out themselves. [191] conduct a lab experiment demonstrating that individuals strategically hide their sexual orientation in anticipation of discrimination in prosocial behavior, a result consistent with that of [189]. Our study complements these studies by focusing on responses to revealing sexual minority status, potentially providing a rationale for the observed underrepresentation of individuals with sexual minority status in many areas.
@@
Second, our findings relate to the broader literature on discrimination. Early work by [138] paved the way for an extensive literature investigating discrimination instances, mostly based on gender and ethnicity, across diverse economic domains [[, see, e.g.,]]arnold2022measuring, kuhn2013gender. A substantial portion of the empirical evidence stems from field experiments, in which researchers use audit and correspondence studies to isolate the causal impact of identity on behavior [[, see, e.g.,]]ayres1995race, oreopoulos2011.666 For recent surveys, see [162] and [168]. [190] provides a detailed review of traditional statistical and taste-based discrimination models, along with a discussion of recent theories that expand on these models. Our study contributes by exploring sentiments against sexual orientation minorities. The concealable nature of sexual orientation allows individuals to anticipate discrimination and strategically choose to hide their identity [191]. This complicates the use of traditional methods to investigate discrimination against individuals based on their sexual orientation and necessitates innovative strategies for understanding sentiments against groups with stigmatized identities, particularly in environments prone to discrimination.
Third, our study contributes to the literature on video games and social interaction platforms [199]. The recognition of video game data’s potential for research is increasing among economists, who are already capitalizing on the abundance and quality of the available data.777 For example, [169] explore the impact of diversity on team performance using data from an online video game, and [195] investigate the effects of artificial intelligence on collaborative production dynamics in controlled laboratory settings using a team-based video game. [144] utilize a self-designed video game to study racial discrimination. However, despite its industry size and the significant fraction of time spent on playing video games, the scientific literature on these social platforms is only slowly increasing.888 [154] highlight that, according to the American Time Use Survey, almost 10 % of leisure time is spent on playing games and computer use. Global revenues of the video game industry were around \$180 billion in $2020$. [196] investigate consumer behavior in the video game industry. They examine the impact of marking products with a gay label on consumers’ demand, finding a significant, albeit short-lived, decrease in demand following the introduction of the gay label. [200] study entertainment preferences in the context of video games and test economic theories of belief-based utility using data from League of Legends. Our study complements this line of research by expanding the understanding of individuals’ behavior within these virtual environments.
@@ 2.1 League of Legends
League of Legends is a globally prominent multiplayer online game developed and published by Riot Games. Originally launched on October $27^{th}$, $2009$, the game attracted a vast player base, averaging $180$ million monthly active players as of $2022$ and peaking at $14$ million players in a single day. It has also achieved substantial financial success, generating \$1.8\\$1.8 billion in revenue in that same year.999 See, e.g., https://prioridata.com/data/league-of-legends.
League of Legends demands significant time and cognitive engagement from its user base. In our sample alone—which follows a subset of players over a seven-month period—users collectively spent 647,966647{,}966 hours in matches. Valuing this gameplay time at the 2022 U.S. federal minimum wage of \$7.25\\$7.25 per hour implies an aggregate opportunity cost of \$4.7\\$4.7 million. This figure represents a conservative lower bound of time spent on the game, as it excludes substantial time spent on pre- and post-match activities.
@@
Before a match begins, players must select a playable character to control during the match. In our analysis, we measure players’ revealed preferences for a specific character by quantifying how frequently they select that character for their matches. Our objective is to investigate whether these preferences undergo any shifts following the disclosure of the character’s sexual orientation. Thus, we devote the rest of this section to exploring the design of characters in League of Legends and the process through which players select their characters for matches.
Each character has a unique set of skills and abilities and is specifically designed to excel in one or two of the distinct roles that players can take on within the team.101010 These roles are labeled top, jungle, mid, bottom, and support. Additionally, characters are crafted with a rich background that adds a narrative dimension to the game but does not have any impact on the game’s mechanics. This is achieved through the creation of detailed biographies and short stories that provide players with a deeper understanding of the character’s history and motivations, thus offering players the opportunity to connect with their chosen characters on a more personal level.111111 The list of all characters along with details about their abilities and histories is available at https://www.leagueoflegends.com/en-gb/champions/.
@@ 2.2 Coming-Out Event
Every year in June, LGBT Pride Month, a dedicated time to honor and celebrate the LGBT community, takes place. Since $2018$, Riot Games has actively participated in this month-long celebration by integrating new content into League of Legends during the month of June. This includes the introduction of in-game cosmetics, such as character skins, as well as emotes that allow players to express themselves in the game. It is important to note that while these additions enhance the visual and expressive elements of the game, they do not alter the game’s mechanics or the characteristics and abilities of the League of Legends characters.121212 We check this in Section 5.1, where we demonstrate that characters’ strength was unaffected by LGBT Pride Month.
Riot has also supported the representation of the LGBT community in the game by unveiling the sexual minority status of specific characters. In our paper, we primarily focus on the character Graves, chosen due to the existence of a well-defined announcement regarding his sexual orientation.131313 For instance, in $2018$, the character Neeko was confirmed as lesbian, marking her as the first openly LGBT character in the game. Furthermore, in $2021$, the characters Diana and Leona were revealed to be lesbians, while the character Nami was declared to be lesbian and polyamorous.
At the beginning of the $2022$ LGBT Pride Month, Riot Games released a short story featuring Graves and Twisted Fate. This story officially unveils Graves’ sexual orientation, establishing him as a gay character.141414 The story unveiling Graves’ sexual orientation also subtly hints at Twisted Fate’s pansexuality, although this is not explicitly stated. We investigate whether this implied revelation has captured the players’ attention in Appendix C. We highlight the relatively low attention directed towards Twisted Fate from players, who were primarily focused on Graves and the explicit establishment of his sexual orientation. As a result, we concentrate our analysis on Graves and his disclosure for a more credible identification of the effects of coming out. The following quotes provide two pivotal passages of the narrative:151515 The whole story is available at https://universe.leagueoflegends.com/en_SG/story/the-boys-and-bombolini/.
@@
This coming-out event closely approximates an ideal experiment where individuals randomly disclose their sexual minority status, thus providing a unique setting to identify the effects of coming out on players’ preferences for Graves.161616 It is crucial to distinguish between the coming-out event and the disclosure of Graves’ sexual orientation. The coming-out event encompasses both Graves’ disclosure and the start of LGBT Pride Month. While this is not a concern for identification, it requires careful interpretation of the findings. To maintain clarity, we generally refer to the effects of the coming-out event in our analysis. Further discussion on this topic is deferred to Section 5.5 and Appendix D.
To ensure the credibility of our identification, it is crucial that the disclosure was not anticipated by players. Figure 1 displays the Google search interest for the query “Graves gay.” We observe minimal interest in this search term throughout the year $2022$, with a remarkable spike occurring during the week of the coming-out event. This pattern supports our assumption of no anticipation and strengthens the credibility of our identification strategy. What is particularly interesting is that this surge in interest surpasses the level observed during the $2022$ League of Legends World Championship (held from September $29^{th}$ to November $5^{th}$), despite Graves being among the top-eight most played characters during the tournament. This finding emphasizes the substantial impact and attention that the coming-out event received from players.
Figure 1: Google search interest over time for the query “Graves gay.” The dashed vertical line denotes the week of disclosure, and the shaded area highlights the League of Legends World Championship.